# Exploring the Limits of Vision-Language-Action Manipulations in Cross-task Generalization

### v2

@@ 6 Conclusions and Discussions
Conclusions. In this work, we have introduced AGNOSTOS, a benchmark for evaluating cross-task zero-shot generalization in robotic manipulation. With 23 unseen tasks across two difficulty levels, AGNOSTOS provides a rigorous testbed for assessing the limits of Vision-Language-Action (VLA) models. Our evaluation of diverse VLA models reveals their significant limitations in unseen task generalization. To address this, we propose X-ICM, a cross-task in-context manipulation method that leverages the cross-task generalization capabilities of LLMs. The X-ICM achieves substantial improvements over existing VLA models, demonstrating robust cross-task zero-shot generalization.
Discussions. While X-ICM significantly enhances cross-task zero-shot generalization, its performance on many unseen tasks, particularly those with both novel objects and motion primitives, remains limited due to LLMs’ challenges in extrapolating beyond pre-training data and in-context demonstrations. In addition, the use of visual information is limited to textualizing object information, which may ignore important visual context in the raw data. To address these limitations, future work could integrate multi-modal reasoning, combining vision, language, and action data to improve generalization to novel semantics. Additionally, leveraging generalizable concepts like object trajectories as a bridge between MLLMs’ perception and dynamic action sequence prediction could reduce cross-task generalization complexity. Scaling X-ICM to diverse robotic embodiments would further enhance its versatility across varied platforms and sensor configurations. We believe the proposed X-ICM benchmark and X-ICM method will inspire future research in generalizable robotic manipulation, paving the way for robots that can seamlessly adapt to open-world environments.

### v3

@@ 6 Conclusions and Discussions
Conclusions. In this work, we have introduced AGNOSTOS, a benchmark for evaluating zero-shot cross-task generalization in robotic manipulation. With 23 unseen tasks across two difficulty levels, AGNOSTOS provides a rigorous testbed for assessing the limits of Vision-Language-Action (VLA) models. Our evaluation of diverse VLA models reveals their significant limitations in unseen task generalization. To address this, we propose X-ICM, a cross-task in-context manipulation method that leverages the cross-task generalization capabilities of LLMs. The X-ICM achieves substantial improvements over existing VLA models, demonstrating robust zero-shot cross-task generalization.
Discussions. While X-ICM significantly enhances zero-shot cross-task generalization, its performance on many unseen tasks, particularly those with both novel objects and motion primitives, remains limited due to LLMs’ challenges in extrapolating beyond pre-training data and in-context demonstrations. In addition, the use of visual information is limited to textualizing object information, which may ignore important visual context in the raw data. To address these limitations, future work could integrate multi-modal reasoning, combining vision, language, and action data to improve generalization to novel semantics. Additionally, leveraging generalizable concepts like object trajectories as a bridge between MLLMs’ perception and dynamic action sequence prediction could reduce cross-task generalization complexity. Scaling X-ICM to diverse robotic embodiments would further enhance its versatility across varied platforms and sensor configurations. We believe the proposed X-ICM benchmark and X-ICM method will inspire future research in generalizable robotic manipulation, paving the way for robots that can seamlessly adapt to open-world environments.
Acknowledgements. This work was supported by the National Natural Science Foundation of China (No. 62306257), the Guangzhou Municipal Science and Technology Project (No. 2024A03J0619 and No. 2024A04J4390) and the Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2023A03J0008), Education Bureau of Guangzhou Municipality.